# LED: LLM Enhanced Open-Vocabulary Object Detection without Human Curated Data Generation

### v5

@@ 1 Introduction
(1) We introduce a lightweight adapter that directly fuses frozen MLLM hidden representations into any object detector, establishing a unified knowledge-injection pipeline and eliminating the need for expensive data synthesis.
@@ 4.2.3 Different LLMs
We evaluated our model on RefCOCO/+/g [40], and the results are shown in Table 5. Combining the evaluation results from OmniLabel, we found that: L ED  +Qwen2-0.5B shows comparable performance to L ED  +InternLM2.5-7B despite $14\times$ more parameters. This suggests a pattern of diminishing returns in scaling LLMs for visual grounding, supporting the hypothesis from [7] that small models can suffice for certain vision–language mappings.
@@ A.3 Experiment Configuration
The hyper-parameters for Stages 1–3 are summarised in 7(a). To investigate the impact of model scale, we vary LLM parameter sizes as detailed in 7(b), with the results reported in Section 4.2.3. 7(c) lists the settings used to evaluate alternative adapter architectures; the corresponding OmniLabel scores appear in Table 3. Finally, 7(d) specifies the setup for examining different decoder layers, and the ablation results are presented in Section 4.4.

### v6

@@ 1 Introduction
(1) We introduce a lightweight adapter that directly fuses frozen MLLM hidden representations into any object detector, establishing a unified knowledge‐injection pipeline and eliminating the need for expensive data synthesis.
@@ 4.2.3 Different LLMs
We evaluated our model on RefCOCO, RefCOCO+, and RefCOCOg [40]; the results are reported in Table 5. Together with the OmniLabel evaluations, we observe that L ED  + Qwen2-0.5B performs on par with L ED  + InternLM2.5-7B despite the latter having $14\times$ more parameters, indicating diminishing returns from scaling LLMs for visual grounding and echoing the hypothesis of [7] that small models can suffice for certain vision–language mappings. In Table 7, we further include D3 and OVDEval benchmarks. Integrating LED-using only the first two layers of Qwen2-0.5B-consistently improves D3 (Full/Pres/Abs) and strengthens OVDEval attributes and relations, with the largest gains in the relationship category. Overall, these results confirm that LED mainly enhances text-conditioned grounding and compositional reasoning while preserving detector efficiency.
@@
Method D3 (AP) OVDEval (AP by sub-category)
Full Pres Abs Color Material Negation Position Relationship
G-DINO 20.7 20.1 22.5 8.9 9.9 23.8 23.8 27.2
G-DINO + LED 21.8 21.7 23.6 11.1 10.1 23.8 23.7 30.9
Table 7: Combined results on D3 and OVDEval Sub-categories. G‑DINO-T’s pre‑training (on O365,GoldG,Cap4M) excludes OVDEval categories “celebrity” (e.g., Taylor Swift), “landmark” (e.g., the Stature of Liberty), and “logo” (e.g., McDonald’s logo), so its AP is almost 0% there and omitted for clarity.
@@ 4.4 Deeper Decoder Layers Offer No Additional Benefit
4.6 Detector with L ED  vs. Large VLMs
We benchmark the InternVL2 family on OmniLabel and D3 and contrast them with lightweight detectors. While InternVL2-8B and InternVL2-76B exhibit strong semantic capacity, they require tens to hundreds of times more parameters and GFLOPs than GLIP-T and G-DINO. By comparison, LED, built on the first two layers of Qwen2-0.5B, adds only 8.7% FLOPs to the detector while providing effective semantic fusion. As Table 8, LED markedly improves G-DINO/GLIP-T at a tiny computational cost, whereas scaling to very large MLLMs is far less efficient for the detection setting. In short, MLLMs contribute generality and rich semantics but are not optimized for efficiency or detection; classic detectors are efficient but semantically shallow. LED bridges the gap by injecting LLM semantics without sacrificing detector efficiency.
Model Params GFLOPs AP categ AP descr
OmDet 610M 640G 13.3% 19.5%
Qwen2.5-VL-3B 3.0B 5.44T 4.2% 4.8%
InternVL2-8B 8.1B 8.47T 4.5% 5.2%
InternVL2-76B 75B 17.92T 15.3% 19.3%
GLIP-T 232M 434G 16.4% 25.8%
GLIP + LED 290M 472G 24.7% 36.8%
G-DINO 172M 412G 16.1% 24.6%
G-DINO + LED 230M 448G 23.7% 36.2%
Table 8: Comparison with SOTA detectors and large VLMs using LED. LED adds a small overhead (first two layers of Qwen2-0.5B; $\sim$8.7% FLOPs) yet yields strong gains, especially on description-style queries.
@@ A.3 Experiment Configuration
The hyper-parameters for Stages 1–3 are summarised in 9(a). To investigate the impact of model scale, we vary LLM parameter sizes as detailed in 9(b), with the results reported in Section 4.2.3. 9(c) lists the settings used to evaluate alternative adapter architectures; the corresponding OmniLabel scores appear in Table 3. Finally, 9(d) specifies the setup for examining different decoder layers, and the ablation results are presented in Section 4.4.